### v1

# Limitations of Existing Work, and Guidelines for Future Research

Abstract
Despite location being increasingly used in decision-making systems employed in many sensitive domains such as mortgages and insurance, astonishingly little attention has been paid to unfairness that may seep in due to the correlation of location with characteristics considered protected under anti-discrimination law, such as race or national origin. This position paper argues for the urgent need to consider fairness with respect to location, termed spatial fairness, by outlining the harms that continue to be perpetuated due to location’s correlation with protected characteristics. This interdisciplinary work connects knowledge from fields such as public policy, economic development, and geography to highlight how fair-AI research currently falls short of correcting for spatial biases, and does not consider challenges unique to spatial data. Furthermore, we identify limitations of the handful of spatial fairness work proposed so far, and finally, detail guidelines for future research so subsequent work may avoid such issues and help correct spatial biases.
fairness machine learning, fair-AI, machine learning, ICML
1 Introduction
The adoption of data-driven decision-making systems has skyrocketed across the board in the last two decades due to their ease of deployment and proficiency in learning patterns from enormous amounts of historical data to make decisions, reducing human effort while increasing efficiency. As the fairness in artificial intelligence (fair-AI) literature (Mehrabi et al., 2021; Pessach & Shmueli, 2022) has noted, however, such decision-making systems can also replicate undesirable trends in historical data, leading to undesirable patterns such as historic racial or gender discrimination being perpetuated further, or worse–possibly even being amplified.
This observation has led to extensive research on developing fairer AI systems, which has centered on protecting against discrimination with respect to personal characteristics or traits protected by anti-discrimination law, commonly referred to in the literature as (legally) protected characteristics, traits, or attributes. For example, the Fair Housing Act of 1968 in the U.S. protects against discrimination based on race, religion, sex, disability, national origin, and familial status in the housing domain (Mehrabi et al., 2021; Lee & Floridi, 2021). There is similar legislation in the U.S. and other parts of the world (e.g., E.U.) protecting against discrimination on the basis of legally protected characteristics in multiple domains, particularly employment, credit, and housing (Xiang & Raji, 2019). However, this singular focus in the computer science community on legally protected characteristics ignores the possibility that other, seemingly neutral characteristics may be correlated to them. For example, research has documented discrimination against racial minorities in the mortgage industry (Pager & Shepherd, 2008; Ross et al., 2008; Massey et al., 2016), despite the Equal Credit Opportunity Act prohibiting creditors from discriminating against applicants on the basis of race (Mehrabi et al., 2021). This discrimination likely occurs not due to an intent to discriminate but due to the presence of the extensive correlation of race with zip codes in the United States due to prejudiced practices such as redlining prevalent in the past (Schill & Wachter, 1994; Woods, 2012).
As such examples show, simply not considering protected characteristics in decision-making, an approach known as ‘fairness through unawareness’ (Grgic-Hlaca et al., 2016; Kusner et al., 2017) is insufficient in practice (Lee & Floridi, 2021). Other fair-AI work has argued to “repair” or fix attributes that may act as proxies for legally protected characteristics (He et al., 2020; Kamiran et al., 2010; Berk et al., 2017; Zafar et al., 2017). However, subsequent research has shown this is ineffective in practice (Corbett-Davies & Goel, 2018; Lee & Floridi, 2021). Other work argues for dropping both the proxies and protected characteristics from the data altogether (Datta et al., 2017), but this is not ideal since a proxy attribute can carry information relevant to decision-making, such as crime rates in zip codes for mortgage decisions.
While many attributes may be correlated with a legally protected attribute, the use of location is particularly troubling due to its potential to be correlated with multiple legally protected characteristics, such as race or national origin (Marshall, 1974; Willborn, 1984; Romei & Ruggieri, 2014; Becker, 2010). Table 1 outlines legally protected characteristics observed to have significant correlation with location, or can be inferred from location data. Extensive research in public policy, economic development, and sociology has documented the extensive spatial segregation that still adversely affects society today. The danger that such historical biases that predominantly disadvantage racial and ethnic minorities (Rothstein, 2015; Schill & Wachter, 1994; Rothstein, 2017) may seep into outcomes of decision-making systems is alarming and warrants greater attention.
Fairness concerning location is further complicated by challenges specific to spatial data, limiting the generalizability of current fair-AI methods, motivating the development of fairness definitions and methods tailored for spatial fairness.
This work advocates for the need for a focus on spatial fairness within fair-AI. We use the interplay of legislation, historical discrimination, and AI systems within the U.S. to motivate greater attention towards bias due to location throughout the paper, however, the issue of discrimination due to location because of its correlation with other attributes is not limited to the U.S. alone (Galster & Friedrichs, 2015; Korsu & Wenglenski, 2010). By drawing on literature in the law, sociology, and economic development as well as computer science, we aim to ensure that future research by computer scientists in fair-AI does not fall into the trap of not being implementable in practice due to its noncompliance with legal or other standards (Xiang & Raji, 2019; Barocas & Selbst, 2016; Kim, 2022). We further delineate challenges specific to location that future work would need to be aware of, and finally draw on relevant research in fields such as public policy and the law to define guidelines for subsequent work in spatial fairness. Our main contributions are:
  • Highlight the importance of systematically addressing spatial fairness within the context of fair-AI, drawing upon scholarly literature in the domains of law and public policy
  • Identify limitations of the handful of spatial fairness definitions and techniques proposed in the computer science literature so far and why they are insufficient
  • Draw from fields of public policy, law, transport geography and economic development to delineate guidelines for future work on spatial fairness
  • Motivate interdisciplinary work that comprehensively brings together relevant knowledge from different domains to push fair-AI forward and address a blind spot in current fair-AI research
The paper is organized as follows. Section 2 motivates spatial fairness by disentangling the complex background on spatial biases and its ongoing societal impact, using growing legal efforts to establish protection against discrimination based on zip codes as a case study. Challenges unique to spatial data that limit generalizability of current fair-AI work to spatial contexts are detailed in Section 3. Section 4 outlines limitations of the handful of existing spatial fairness definitions and techniques. Guidelines for future spatial fairness work to avoid such pitfalls and comply with legal and public policy standards are delineated in Section 5. We conclude with Section 6.
Attribute Protected at level Correlation References & Notes
w/ location
[…]
Protected in housing in 15 states (pol, )
[…]
Table 1: Protected characteristics in housing (federal: Fair Housing Act; state: Policy Surveillance Program), credit (Equal Credit Opportunity Act, and employment (Equal Employment Opportunity Commission. (Mehrabi et al., 2021; Chen et al., 2019; pol, )
2 Context and Motivation: Location bias
We discuss the history of spatial segregation within the U.S., resulting in pervasive location-based bias and discrimination even today, and outline the case for legal protection for discrimination on the basis of location, particularly zipcodes.
2.1 Spatial unfairness in the real world
Although the fair-AI literature has yet to explore bias due to location in depth, it has been extensively covered in fields such as public policy, urban planning, and the law (Schill & Wachter, 1994; Rothstein, 2017; Baum et al., 2015). For example, neighborhoods in the United States have been historically correlated with race, a legally protected characteristic in many domains (e.g. housing, employment, credit), in complex ways due to prejudiced practices such as redlining [^1] Redlining is the discriminatory practice of systematically denying individuals access to mortgages, insurance, and other financial services living in certain neighborhoods, based on their race or ethnicity. (Woods, 2012; Rothstein, 2017). Even though the practice of redlining was prohibited by the Fair Housing Act in 1968 [42 U.S.C. §§ 3601 - 3619], unfortunately, its effects continue to linger (Angwin et al., 2017; Mehrabi et al., 2021). Zip codes in many parts of the country still correlate with race, potentially acting as a proxy for race in data-driven decision-making scenarios where zip codes are considered as a feature, such as in mortgages and insurance. This has been frequently documented. For example, the non-profit ProPublica showed that drivers who live in (racial) minority neighborhoods are charged higher insurance premiums than drivers of same risk levels who reside in predominantly white neighborhoods (Angwin et al., 2017). Research has also found that prices for ride-hailing trips in the city of Chicago from or to neighborhoods with larger non-white populations or higher poverty levels were significantly likely to be charged higher prices (Pandey & Caliskan, 2021; Saxena et al., 2023b).
Neighborhoods (and their zip codes) may also be correlated with immigrants, i.e., neighborhoods where immigrants are more likely to live, which can act as a proxy for national origin, another characteristic protected against discrimination in domains such as housing, credit, and employment under U.S. law (Mehrabi et al., 2021). Furthermore, neighborhoods may also be correlated with income (Mode et al., 2016). In addition to the ride-hailing example mentioned above which found higher trip prices significantly associated with neighborhoods with higher poverty levels, residents of low-income neighborhoods may also be subjected to other forms of income-related discrimination. For example, families receiving housing vouchers often face barriers in moving to lower-poverty and higher-income neighborhoods due to discrimination by landlords against their source of income (Mazzara & Knudsen, 2019). While not illegal at the federal level, some states have laws protecting people from discrimination based on source of income (Table 1). As a result, many families receiving housing vouchers still end up in low-income, racially segregated neighborhoods despite being able to afford homes in better neighborhoods (Rothstein, 2015). Such families might suffer doubly if they are also then subjected to spatial discrimination due to their low-income, racially segregated neighborhood.
Despite racial segregation in housing being banned in the U.S. in 1968, spatial segregation by race has persisted, resulting in “racially homogenous public institutions that are geographically defined,” (Fulwood III, 2017). As a consequence, such neighborhoods have reduced access to economic, educational, and social opportunities (Chetty et al., 2016; Li et al., 2013), which further hinders future prospects for residents in those neighborhoods. This also manifests in house prices, with a study finding that the racial composition of a neighborhood was a stronger determinant of a home’s appraised value in 2015 than in 1980 (Howell & Korver-Glenn, 2021). Furthermore, research has shown that prices for houses in majority-white neighborhoods are significantly higher than for similar houses in neighborhoods that happen to be majority-black, worsening the wealth gap between whites and racial minorities, especially African Americans, since a home is typically a family’s most significant financial asset according to a report by Pew Research Center (Taylor et al., 2011). There has been no significant change over the past few decades in the manner in which majority-black neighborhoods continue to be encircled by spatial disadvantage (Sharkey, 2014). Ethnographic research has documented that even middle class African-American families with an annual income of at least \$100,000 tend to live in neighborhoods with the same disadvantages as the average white family with an annual income of \$30,000 or lower (Sharkey, 2014).
Due to such extensive correlation of spatial data with legally protected attributes, decision-making systems that have been developed to be fair with respect to legally protected characteristics directly may still exhibit unfair behavior.
While we focus on anti-discrimination law in the United States to motivate this problem, unfortunately such spatial inequities are not restricted to the U.S. alone. Research in fields such as economics and public policy have documented rising spatial inequities in countries across the world (Kanbur & Venables, 2005; Kochendörfer-Lucius & Pleskovic, 2009; Zhao & Tong, 2000; Feldman & Storper, 2018).
2.2 Legal support for spatial fairness
The law in the United States recognizes two forms of discrimination for protected characteristics or classes (Lee & Floridi, 2021; Baum et al., 2015): disparate treatment and disparate impact. In layman terms, disparate treatment can be thought of as direct or intentional discrimination while disparate impact can be viewed as indirect or unintentional discrimination (Barocas & Selbst, 2016; Primus, 2010). While disparate treatment focuses on the intent to discriminate, disparate impact focuses on outcomes that are discriminatory towards members of a protected class despite a seemingly neutral decision-making policy.
For both forms of discrimination, U.S. law provides protections against discrimination on the basis of legally protected characteristics or attributes in domains such as housing [42 U.S.C. § 3601 et seq], credit [15 U.S.C. 1691 et seq.], and employment [42 U.S.C. § 2000e]. Some states define additional protected attributes, for example, marital status is also a protected class in the state of California in the domain of employment (Gelb & Frankfurt, 1982) although it is a protected characteristic only in the domain of credit at the federal level. Typically under U.S. anti-discrimination law, protected characteristics are those which are considered immutable, i.e., characteristics that were not chosen and cannot be changed over time (Hoffman, 2010), such as race, color, and sex. Race and color were among the first classes to be considered immutable in the eyes of the law, and deemed protected classes. In the recent past, however, this definition has been expanded in a series of court judgements to include characteristics that are deemed difficult or too significant to ask someone to change. For instance, in June 2020, the U.S. Supreme Court issued a landmark judgement that found that discriminating against a person on the basis of their sexual orientation or gender identity in employment violates the protection for discrimination under sex in Title VII of the Civil Rights Act of 1964 [Bostock v. Clayton Cty., 140 S. Ct. 1731 (2020)].
We argue that residential zip codes can also be considered an immutable characteristic for a section of the population, thereby deserving protection against discrimination on the basis of zip codes. Common sense dictates that it is not easy or straightforward for an individual or family to move to a lower-poverty or less segregated neighborhood in order to access better economic or educational opportunities. Even when disadvantaged families can afford to live in higher-income, more integrated neighborhoods with the support of housing vouchers they still cannot move due to source of income discrimination by landlords (Mazzara & Knudsen, 2019). Research has also shown that such discrimination overwhelmingly affects minorities, perpetuating systemic racism “and denies renters of color, especially Black renters, equal access to housing opportunities.” (Coa, 2023) Therefore, for low socioeconomic status families, where they live can be considered an immutable characteristic – they simply do not have the ability to move even if they wish or can afford to. Furthermore, they are forced to remain in their low-income neighborhoods which are likely to be correlated with race due to historical segregation, suffering the negative consequences associated with that in addition to the discrimination due to their source of income.
Due to the amalgamation of these negative factors, the groups affected most severely are people of color, people with disability, and women, which may be considered disparate impact. This line of thinking is gaining traction. After ProPublica published analyses documenting higher insurance premiums for drivers living in minority neighborhoods than similar-risk drivers in majority-white neighborhoods (Angwin et al., 2017), former Illinois State Senator Jacqueline Collins proposed a bill banning the use of zip codes in determining insurance quotes as they can act as proxies for race (Salcedo, 2017). While that did not pass, Illinois State Representative Will Guzzardi introduced the House Bill 2203 in February 2023 that proposes to end the use of non-driving factors including zip codes in setting insurance rates due to their well-known correlation with race.
While the fate of House Bill 2203 remains to be seen, the tangible impact on society of location’s correlation with protected attributes cannot be denied. Regardless of legal protections afforded to location bias, the fair-AI community has an ethical responsibility to strive to address it given the extensive impact our systems have on society. In this goal, computer scientists must contend with challenges specific to spatial data in addition to bias, as we shall see next.
3 Problems unique to spatial data
Much fair-AI work cannot be directly extended to spatial fairness since it does not account for certain challenges specific to analyzing spatial data. We detail them below.
3.1 Dimensionality
Geographic locations are two-dimensional, considerably increasing complexity (McMillen, 2010). Furthermore, the choice of level of detail for spatial data is not straightforward. For example, a geographic area may be represented at the level of houses, blocks, neighborhoods, or census tracts. While level of detail may be application-dependent, a finer level of detail increases computational complexity (Haining, 2003). Dimensionality reduction can be tricky with big data, and fairness concerns add another layer of complexity by making it a multi-objective problem (i.e., reducing dimensionality while satisfying fairness constraints). Fair dimensionality reduction, a nascent area, is non-trivial, and an open problem. For example, Samadi et al. (2018) show PCA often has higher reconstruction error on one subgroup than another, e.g., women versus men, and this may occur even when the data contains a similar number of samples from both groups. They propose fair-PCA (Samadi et al., 2018), but it requires $d+1$ dimensions for 2 protected subgroups, and is prohibitively slow in practice, while Tantipongpipat et al. (2019) propose a polynomial-time Fair-PCA algorithm for $k=2$ subgroups. Spatial data’s correlations with many protected characteristics (e.g., race, ethnicity), with each characteristic having multiple subgroups, complicates fair dimensionality reduction in spatial decision-making. Techniques in other areas such as r-trees (Guttman, 1984) that are designed to take advantage of the unique relationships between the two spatial dimensions to improve performance for multi-dimensional spatial data can offer insights for leveraging unique spatial dynamics to improve upon generic dimensionality reduction methods for spatial scenarios.
3.2 Complexity of computing spatial network distance
Computing distance between two locations involves crucial design choices with practical implications. The commonly used Euclidean distance, i.e., the distance between points $A$ and $B$ in a straight line, is straightforward to compute but does not reflect road network distance, i.e., the distance between $A$ and $B$ by road. More complex to compute, road network distance often differs from Euclidean distance in practice (Sander et al., 2010; Apparicio et al., 2003). Moreover, in certain decision-making scenarios considering travel time, i.e., the time it takes to travel between $A$ and $B$, is crucial. For example, houses equidistant to a hospital in road distance may still have very different reachability to the hospital in practice due to traffic patterns and/or lack of public transport making the travel time for houses in one zip code much worse than another (Anastasiou et al., 2019). The time-dependent, dynamic nature of traffic adds further complexity to the problem (Li et al., 2017). Furthermore, as distance metrics are central to individual fairness, increased complexity due to road network distance and travel time increases difficulty for individual spatial fairness as well.
3.3 Continuity of space
Geographic spaces are continuous in nature, and not as easy to analyze as discrete spaces (Xie et al., 2022). Discretizing spatial data is not straightforward due to features of spatial data such as the spatial structure, association, and heterogeneity (Cao et al., 2014; Haining, 2003). Furthermore, geographical partitionings are heavily context-dependent, and less likely to have partitionings that are universally well-accepted than other continuous variables such as income which has well-defined thresholds (e.g., for income tax). Additionally, when the decision-making scenario is spatio-temporal, i.e., involves travel time, there is one more continuous dimension – time, further complicating matters. Finally, achieving fairness with respect to continuous attributes is non-trivial. Most fair-AI work focuses on discrete variables (e.g., gender: male/female) (Mehrabi et al., 2021). Although recent work explores fairness with continuous attributes, many depend on computationally intractable statistical independence measures in practice (Jiang et al., 2022; Mary et al., 2019; Jha et al., 2021; Creager et al., 2019), lack estimation accuracy guarantees (Jiang et al., 2022; Cho et al., 2020; Roh et al., 2020), or are not sensitive to protected subgroups that are less likely to occur in the population (Jiang et al., 2022). The combined complexities of continuous in two- or three- (with time) dimensions spatial data and achieving fairness with respect to continuous attributes make this a non-trivial task.
3.4 Modifiable Areal Unit Problem
The modifiable areal unit problem (MAUP), a phenomena observed in geography, is a form of statistical bias that arises when point-based estimates of spatial phenomena are aggregated into spatial partitions; the results from such an analysis will be influenced by the shape and scale of the spatial partition (Wong, 2004; Fotheringham & Wong, 1991). For a given space, the manner in which we define geographic boundaries or vary the level of geographic partitionings may change the perception of fairness. In other words, data collected for the same region but at different spatial scales will give inconsistent or different results. A common example of MAUP in the real-world is gerrymandering, which we discuss in Section 5.
3.5 Spatial autocorrelation
“Everything is related to everything else, but near things are more related than distant things.”
– Waldo Tobler, the First Law of Geography Finally, there is the challenge of spatial autocorrelation. The First Law of Geography states that nearby things are more related than distant ones, violating the independence assumption in standard statistical tests, requiring careful selection of statistical methods. Recent works in ML study such spatial dependence in different fields. Deppner et al. (2022) propose a spatial cross-validation strategy to reduce spatial bias in tree-based algorithms commonly used for price and rent modeling in the real estate industry. In the geosciences, work explores spatial autocorrelation when spatial features like spatial lag are used in random forests (Liu et al., 2022), and incorporating spatial autocorrelation measures in ML modelling for predicting geotechnical information, such as ground surface elevation (Kim et al., 2023), and increasing the prediction accuracy of heavy metals in soil (Sergeev et al., 2019). However, none of these works look at spatial fairness or spatial correlation with respect to legally protected attributes. Recently, Majumdar et al. (2022) propose a hypothesis testing methodology to detect the presence of spatial autocorrelation and its strength in presence of correlation with a protected attribute, although they do not propose techniques to mitigate the observed spatial bias. Thus, despite significant practical applications, mitigating spatial autocorrelation and its interplay with protected attributes remains a deeply understudied area.
These complexities of spatial data hinder use of fair-AI work in spatial scenarios, motivating dedicated spatial fairness, as recent works have limitations hindering their adoption.
4 Limitations of current spatial fairness work
Initial attempts at spatial fairness definitions and techniques fall short in significant ways, as we detail below.
4.1 Spatial fairness definitions: Legal soundness
Many fairness definitions have been proposed in fair-AI research (Verma & Rubin, 2018; Mehrabi et al., 2021), but most are not directly applicable to spatial fairness due to the challenges specific to spatial data noted in Section 3. Moreover, many fairness definitions proposed so far will not hold up to legal scrutiny (Xiang, 2020; Wachter et al., 2021). They may not be deemed legal in a court of law, and may thus be inapplicable in practice. Worryingly, many definitions may be incompatible with U.S. anti-discrimination as well as E.U. non-discrimination law.
Incompatibility with anti-discrimination law may also be an issue for the handful of spatial fairness definitions proposed so far. For example, (Shaham et al., 2022) propose two types of spatial fairness: distance-based spatial fairness and zone-based spatial fairness. The former is concerned with decision-making scenarios based on an individual’s distance from a reference point, such as a shop running a marketing campaign offering discount vouchers. However, it is legal to make decisions based on an individual’s distance from a reference point. The latter definition, zone-based fairness, considers scenarios where an individual’s location is taken relative to specific zones. It is legal to make a decision based on an individual’s location relative to specific zones, unless the zone is directly correlated with a protected attribute, which current work does not distinguish. Indeed, school districts, banking and mortgage industries, et cetera use zones to determine school assignment or in the computation to determine mortgage rates. To ensure adoption in the real-world, it is imperative that future spatial fairness work is in compliance with legal definitions of anti-discrimination.
4.2 Spatial fairness techniques: limitations
4.2.1 Do not close the loop
Perhaps the most concerning limitation of most spatial fairness techniques so far is that current research on spatial fairness does not yet close the loop: They do not demonstrate that their proposed spatial fairness methods addressed unfairness present due to location’s correlation with protected characteristics (Shaham et al., 2022; Sacharidis et al., 2023). Although discrimination on the basis of legally protected characteristics is the motivation for the spatial fairness techniques proposed so far, there are no experiments demonstrating techniques that can reduce discrimination on the basis of such attributes, like race.
The incompatibility of most spatial fairness definitions with U.S. and E.U. laws may hinder their practical adoption. Moreover, these techniques must unequivocally demonstrate improvement in fairness with respect to protected characteristics before being considered for adoption.
4.2.2 Take agency away from people
Some of the formulations of spatial fairness proposed so far use a fuzzy boundary instead of hard boundaries as are typically employed in domains such as education (e.g., school districts), banking, or housing. While this may have some technical advantage, a significant consequence of this would be to take agency away from people. For example, with clear school-district boundaries, a low-income family with school-age children can choose to make sacrifices to live in a better school district (Hristov & Khare, 2016). If such boundaries are fuzzy, however, then this can impact the decision-making ability of individuals and families in a deeply negative manner. Wealthy families will still be able to ensure a good education for their children by living squarely in the heart of good school districts or simply paying for private schools. But lower and middle income families will suffer due to the opaqueness of the fuzzy boundary, and not being able to assess the chance that a particular address may have of securing their children admission in a good school. Research has shown that one of the barriers for disadvantaged families to be able to benefit from public benefits they are eligible for is lack of knowledge of eligibility rules (Wu & Eamon, 2010; Anderson, 2002; Stuber & Kronebusch, 2004; Zedlewski, 2003) and complicated application processes (Wu & Eamon, 2010; Morrow & Horner, 2006; Zedlewski et al., 2002). Therefore, increasing the complexity of assessing which school district an address may belong to will overwhelmingly harm those least likely to be equipped to deal with it efficiently.
4.2.3 Do not build upon existing knowledge
Research is about building on existing work to stand on the shoulders of giants. A considerable amount of work explores different aspects of spatial fairness with respect to protected attributes. For example, literature in spatial scan statistics (Wong, 2004; Fotheringham & Wong, 1991) and transport geography (Guo & Bhat, 2004; Mitra & Buliung, 2012) explore MAUP from different angles, which can offer ideas for how to grapple with MAUP in spatial fairness. Similarly, work in public policy studies spatial segregation patterns with respect to race, ethnicity, and immigration in housing and schools across the country (Schill & Wachter, 1994; Sanchez, 2006; Saiz & Wachter, 2011; Schill & Wachter, 1995). Research in economics can help offer guidance on which spatial measures help improve economic development, especially for the most disadvantaged (Chetty et al., 2016; Brata et al., 2009; Glasmeier, 2018)
5 Guidelines and Future Directions
We define guidelines to help future spatial fairness research avoid the pitfalls and challenges outlined in Sections 3 and 4, and delineate future directions.
5.1 Close the loop
Research exploring spatial fairness so far does not close the loop, i.e., demonstrate that their proposed work reduces negative effects on people belonging to protected classes. Spatial fairness work should depict empirical experimental evaluation of the extent of spatial bias with respect to protected attributes in the problem domain with and without the proposed methodology and its impact on other problem objectives (e.g., prediction accuracy; efficiency of matching students to schools). For example, suppose a methodology places fairness constraints on location $L(x)$ to mitigate its bias associated with protected attribute $x$, getting $L^{\prime}(x)$ in the process. Experiments must demonstrate the extent of bias between $L(x)$ and $L^{\prime}(x)$ with $x$, and the trade-offs (if any) of mitigating bias with accuracy on outcomes, in addition to comparisons with previously proposed methods.
5.2 Avoid disparate impact
Under U.S. laws, practices that unintentionally harm people of a protected characteristic, such as in employment and housing, may be considered discrimination (Corbett-Davies & Goel, 2018; Feldman et al., 2015), or disparate impact in legal terms. Research has shown the potential for computational decision-making models to perpetuate disparate impact (Barocas & Selbst, 2016). To address this, a growing body of work advocates for the development of disparate impact assessments, potentially modeled on similar analyses in industries that must comply with non-discrimination laws (MacCarthy, 2017). Other work also argues for such vigilance or assessment to continue after model deployment, perhaps by external auditors, to ensure models function as intended and have no unpleasant unintended consequences (Reisman et al., 2018; Saxena et al., 2023a). Assessing disparate impact in computational systems is not trivial, with healthy debate in the fair-AI community about how to detect and assess disparate impact. Work has proposed tests and methods to detect and measure disparate impact in data-driven systems (Feldman et al., 2015; Mehrabi et al., 2021; Metcalf et al., 2021; Reisman et al., 2018), and critiqued tests and their misalignment with U.S. anti-discrimination law (Selbst et al., 2019; Xiang & Raji, 2019). Assessing disparate impact in spatial decision-making, however, must incorporate the unique challenges of location and spatial assessments of outcomes with respect to protected groups.
In domains like employment, credit, and housing, tools that measure such disparate impact and are compliant with U.S. and E.U. anti-discrimination laws would likely enjoy great demand by supporting companies in prevention, and empowering individuals to identify if they have been subject to unfair effects of discriminatory practices.
5.3 Location: a protected or immutable characteristic
Immutability is deeply associated with anti-discrimination law in the U.S. (Hoffman, 2010). Recent interpretations are a characteristic “that either is beyond the power of an individual to change,” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985)] or in other terms, “either unchangeable in absolute terms or so fundamental to identity or conscience that individuals effectively cannot and should not be required to change it” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985).] or “changing it would require great physical difficulty, such as requiring a major physical change or a traumatic change of identity.” [Watkins v. U.S. Army, 875 F.2d 699, 726 (9th Cir. 1989)] (Hoffman, 2010), thereby establishing an immutable characteristic need not be unchangeable to be considered immutable under anti-discrimination law. This, in addition to ongoing efforts to afford protection against discrimination on the basis of zip codes, brings us to the next guideline: consider location a protected attribute, or at least an immutable one. We outline example use cases below.
In domains with demonstrated history of significant location-related discrimination, zip codes could be considered a joint protected attribute with the legally protected attribute it perpetuates bias against. For example, an intuitive measure of individual fairness is to treat similar people similarly (Dwork et al., 2012). With location and race as a joint protected attribute, two applicants who are similar to each other on non-protected aspects should receive similar decisions, regardless of their race and location. When impractical, location should be treated as an immutable attribute. For example, some fair- and explainable-AI research aims to provide applicants with actionable points they can improve to increase their likelihood of a positive decision (Ustun et al., 2019). Suggested actions should, however, be reasonably achievable. For example, one can recommend increasing annual salary by \$2,000, but if the salary of an applicant with a PhD degree is low for someone with a PhD degree, one cannot recommend them to undo their PhD. In a similar vein, for low-income individuals and families, residence should be treated as immutable, i.e., something an applicant of a particular socio-economic status cannot reasonably change since they commonly do not have the freedom to make this choice, sometimes even if they can afford to (e.g., due to source of income discrimination (see Sec 2.1)).
5.4 Be resistant to MAUP
MAUP is a dilemma unique to spatial analyses (Sec 3.4). At a high level, two types of spatial biases can distort results when analyzing spatial data. The scale effect occurs when point data is aggregated at different levels of aggregation. Despite using the same point data, different aggregation levels can represent different patterns. On the other hand, the zonal effect occurs when spatial is grouped by different artificial boundaries, leading to different results. Research over the years has found MAUP to exhibit a high level of unpredictability in terms of its intensity and effects in multivariate statistical analyses, significantly surpassing the challenges associated with MAUP in univariate or bivariate statistical analysis (Fotheringham & Wong, 1991)
A classic real-world example that can highlight the dangers of MAUP is the issue of gerrymandering (Nelson & Brewer, 2017; Manley, 2021), the practice of redrawing political boundaries or election districts to serve the interests of a particular political party. There is work in the fields of spatial statistics and geography for methods to address MAUP generally (Tuson et al., 2019; Hennerdal & Nielsen, 2017) or in specific applications such as traffic safety (Xu et al., 2018) or designing zonal districts without falling into the MAUP trap that research on spatial fairness can potentially draw from (Martin, 2003; Tuson, 2022).
5.5 Different definitions for public and private sectors
After extensive debate within the fair-AI community, it is well accepted that different fairness definitions may be necessary for different decision-making contexts (Saxena et al., 2019; Mehrabi et al., 2021). This is likely true of spatial fairness definitions as well, given the diverse criteria and objectives in different decision-making domains. For instance, public policy-driven spatial decisions aim for fairness without profit, such as equitable assignment of students to schools or optimal locations for public services such as hospitals to ensure greater access. In contrast, private companies may seek profit while aiming for spatial fairness. Moreover, industries like mortgage or insurance must adhere to specific state and federal laws prohibiting discrimination.
Hence, articulating the specific domain(s) (public or private) and industries targeted by a proposed spatial definition or method and elucidating the conditions and laws it aims to abide by upon its introduction would facilitate subsequent research endeavors. This includes the evaluation of whether the stipulated conditions and laws are actually adhered to in practice. Since a particular spatial fairness definition or technique may not be generalizable to all spatial decision-making scenarios, doing so would establish a foundation for practical applications and simultaneously ground the work in the possibility that the proposed method is not guaranteed to be generalizable beyond the specified contexts.
5.6 Serve as a tool to measure spatial inequities
Ideally, a spatial fairness definition or methodology should also serve as a tool to detect (or audit for) the presence of spatial inequities. This is not always the case, as Sacharidis et al. (2023) show the measure proposed in Xie et al. (2022) cannot be meaningfully used to audit for spatial fairness or to find which regions are spatially unfair. They go on to propose a methodology that can be utilized to audit for spatial fairness with respect to positive/negative outcomes, but do not consider the spatial regions’ correlation with protected characteristics. Fortunately, there is some work in the social sciences exploring auditing for inequities, including spatial inequities, which may serve as a foundation for spatial fairness and auditing. For example, Saporito and Maliniak (2022) propose a technique to distinguish between legislative districts that are “unnaturally packed” with racial minorities (i.e., are segregated) from districts that are not to aid the study of racial gerrymandering.
Tools for auditing spatial fairness can have significant, meaningful impact in multiple ways. First, by helping computationally implement recent initiatives passed by the justice department to audit lending practices to assess when an individual’s race, national origin or neighborhood had an impact in their access to credit. Second, by helping automate spatial auditing analyses of changes in compositions of redlined neighborhoods over time to support appropriate remedial policy decisions. For example, Perry and Harshbarger (2019) found that housing proposals by some 2016 presidential hopefuls meant to address persisting harms from redlining would not have benefited African Americans as much as intended because the racial makeup of redlined neighborhoods in some areas had changed over time. Spatial fairness tools can make such auditing easier to carry out, helping ensure policies benefit the intended recipients.
5.7 Good faith effort to explore relevant environmental, social, and economic issues in social planning
Fair-AI researchers’ familiarity with public policy and economic development recommendations regarding location (Squires & Kubrin, 2005; Lewis & Hamilton, 2011; Li et al., 2013; Kaplan, 1998; Osypuk & Acevedo-Garcia, 2010) would significantly help integrate their research into practical applications. While it cannot be expected for all computer science research to abide by public policy guidelines, fair-AI researchers should strive to be familiar with public policy and economic development recommendations that are directly relevant to their area of research. Integration of practices recommended to reduce inequities would ensure fair-AI techniques do not unintentionally perpetuate such inequities themselves if implemented in the real world.
6 Conclusion
This work argues for the importance of spatial fairness. We delineate the need to address persisting bias in spatial decision-making contexts and legal efforts proposed to address it, and outline problems unique to spatial analysis that compound the difficulties in addressing spatial bias computationally. Furthermore, we identify drawbacks of the handful of spatial fairness definitions and techniques proposed so far to highlight why they are insufficient. Finally, we establish a set of guidelines to support future spatial fairness research avoid similar pitfalls and achieve maximal impact with adoption in the real world.
Impact Statement
This position paper presents an argument (or position) to bring attention to spatial fairness in the field of machine learning with the goal of motivating future work on mitigating location-related biases in society. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.